OpenStudy (anonymous):

For h is not equal to 0, which of the following is equivalent to A(x+h)^2-Ax^2/h choices are ~ Ah ~ 2Ax+Ah ~Ax+Ah ~ 2Ax Thanks :)

OpenStudy (anonymous):

\[ \frac{A (h+x)^2-A x^2}{h}=\frac{A h^2+2 A h x}{h}=A h+2 A x \]